OpenStudy (anonymous):

Doreen can wash her car in 15 minutes, while her younger brother Dave takes twice as long to do the same job. If they work together, how many minutes will the job take them?

OpenStudy (calculusxy):

What is twice the number of 15?

OpenStudy (calculusxy):

15*2=minute of how long it takes Doreen's brother.

OpenStudy (calculusxy):

So if 15*2=30 Then 15+30=45 minutes altogether

OpenStudy (calculusxy):

If you like my reply, you can give me a medal or be my fan.